### v1

# Scaf-GRPO: Scaffolded Group Relative Policy Optimization for Enhancing LLM Reasoning

Abstract
Reinforcement learning from verifiable rewards has emerged as a powerful technique for enhancing the complex reasoning abilities of Large Language Models (LLMs). However, these methods are fundamentally constrained by the “learning cliff” phenomenon: when faced with problems far beyond their current capabilities, models consistently fail, yielding a persistent zero-reward signal. In policy optimization algorithms like GRPO, this collapses the advantage calculation to zero, rendering these difficult problems invisible to the learning gradient and stalling progress. To overcome this, we introduce Scaf-GRPO (Scaffolded Group Relative Policy Optimization), a progressive training framework that strategically provides minimal guidance only when a model’s independent learning has plateaued. The framework first diagnoses learning stagnation and then intervenes by injecting tiered in-prompt hints, ranging from abstract concepts to concrete steps, enabling the model to construct a valid solution by itself. Extensive experiments on challenging mathematics benchmarks demonstrate Scaf-GRPO’s effectiveness, boosting the pass@1 score of the Qwen2.5-Math-7B model on the AIME24 benchmark by a relative 44.3% over a vanilla GRPO baseline. This result demonstrates our framework provides a robust and effective methodology for unlocking a model’s ability to solve problems previously beyond its reach, a critical step towards extending the frontier of autonomous reasoning in LLM.
Figure 1: Scaf-GRPO overcomes the learning cliff with minimal guidance, outperforming vanilla GRPO grpo and the prefix-based LUFFY yan2025learning across challenging math benchmarks on Qwen2.5-Math-7B. By injecting strategic, hierarchical hints, our method unlocks the model’s potential on difficult problems, achieving superior overall performance.
1  Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in complex reasoning tasks across diverse domains such as mathematics, programming, and logic DeepSeekR1 ; openaio1 ; muennighoff2025s1 ; min2024imitate . A key driver of these advancements is Reinforcement Learning from Verifier Rewards (RLVR) DeepSeekR1 ; simplerlzoo ; liu2025understandingr1zeroliketrainingcritical , a paradigm where models learn to generate sophisticated reasoning paths by exploring diverse strategies and receiving feedback on their final outcomes. This approach eliminates the need for expensive, step-by-step human annotations by rewarding only the final correct answer, enabling models to autonomously discover effective problem-solving procedures DeepSeekR1 .
However, the efficacy of RLVR is severely constrained by a fundamental challenge we formalize as the “learning cliff.” This phenomenon occurs when a model confronts a subset of problems that lie significantly beyond its current capabilities. For these problems, all exploratory attempts consistently fail, leading to two critical and cascading consequences: (1) Reward Signal Loss: The model receives a persistent zero-reward signal for this entire subset of problems. (2) Vanishing Gradients: In algorithms like GRPO grpo , the advantage signal provides the learning gradient. When all rewards are zero, the advantage collapses to zero, providing no gradient for the policy to learn from DAPO .
Consequently, these difficult problems become “invisible” to the policy update. As our empirical analysis in Figure 2 illustrates, these problems form a persistent “long tail” of challenges that the model cannot conquer autonomously. This long tail represents a critical bottleneck, as it prevents the model from leveraging the most difficult examples to achieve a higher level of competence.
To address the learning cliff, a prevailing strategy has emerged: incorporating off-policy guidance from a more capable “teacher” policy yan2025learning ; Prefix-RFT ; StepHint ; BREAD . These methods typically work by providing the student model with a prefix of a correct “golden” solution and tasking it with generating the remainder of the reasoning path. While this ensures a positive reward signal, this prefix-continuation paradigm introduces significant issues. It creates a distributional mismatch between the teacher-generated prefix and the student-generated suffix, necessitating complex algorithmic corrections like policy shaping yan2025learning or hybrid SFT-RL objectives Prefix-RFT that can introduce bias and training instability. More critically, this “on-rails” guidance forces the model down a predetermined path, stifling its ability to explore alternative, potentially more novel or efficient, reasoning strategies.
To address this challenge, we propose Scaf-GRPO (Scaffolded Group Relative Policy Optimization). Our framework is inspired by the pedagogical theory of Scaffolding Scaffolding , a teaching method of providing temporary support that fades as learners improve. We apply this principle by providing hierarchical, minimal, progressive assistance to help the model bridge its capability gaps, rather than enforcing a rigid solution prefix. This in-prompt scaffolding approach is guided by two primary objectives: first, to maintain policy consistency by having the model process both the problem and the hint under a single, unified policy, thereby avoiding the distributional mismatches of prefix-based methods. Second, to preserve exploration flexibility, as our hints act as “signposts” rather than “railroads,” guiding the model without fixing its path and allowing it to discover its own unique solution strategies.
Our framework operates in two carefully designed phases. It first employs a guidance exemption period to distinguish “true-hard” problems from “pseudo-hard” ones that the model can solve on its own with more training. For true-hard problems, it then activates hierarchical hint-guided exploration, providing progressively concrete hints (from abstract concepts to concrete steps) until the model can generate a correct solution. By rewarding the model for succeeding with the most abstract hint possible, Scaf-GRPO encourages the internalization of reasoning skills rather than the memorization of solutions. Our contributions are as follows:
  • We propose Scaf-GRPO, a novel training framework inspired by pedagogical scaffolding addressing the “learning cliff” issue in RLVR. It provides hierarchical, minimal, and progressive guidance via in-prompt hints instead of fixed solution prefixes. This approach maintains policy consistency while preserving the model’s exploratory autonomy, thereby overcoming the key limitations of existing guidance methods.
  • We demonstrate the effectiveness of Scaf-GRPO through extensive experiments on several challenging mathematics benchmarks. On the Qwen2.5-Math-7B model, our method achieves a significant relative improvement of 12.6% over the vanilla GRPO baseline and a 9.2% relative gain over strong prefix-based guidance methods like LUFFY.
  • We demonstrate the broad applicability and robustness of Scaf-GRPO across diverse models. Our experiments show consistent performance gains on different architectures (Qwen, Llama), scales (1.5B to 7B), and specializations (math-tuned, instruction-tuned, and Long-Chain-of-Thought), establishing Scaf-GRPO as a versatile and model-agnostic framework enhancing LLM reasoning.
2  Related Work
Reinforcement learning from verifier reward.
The success of DeepSeek-R1 DeepSeekR1 establishes Reinforcement Learning from Verifier Reward (RLVR) as a paradigm for enhancing the reasoning capabilities of Large Language Models (LLMs). In RLVR, models are trained using feedback from an external verifier that provides an outcome-based reward (e.g., correct/incorrect) for a generated solution. The success of DeepSeek-R1 DeepSeekR1 demonstrates that even with sparse, binary rewards, models can learn reasoning strategies. Subsequent research has built upon this foundation, focusing on enhancing algorithmic stability through debiasing techniques liu2025understandingr1zeroliketrainingcritical ; DAPO , or designing more informative rewards to improve sample efficiency, such as using length penalties to mitigate overthinking L1 ; jin2025recut ; chen2024not or token-level signals to provide denser feedback High-Entropy-token ; Dual-token .
Figure 2: Training dynamics of Qwen2.5-Math-1.5B. (a) Scaf-GRPO overcomes the learning cliff by continuously solving zero-reward problems where vanilla GRPO plateaus. (b) This translates to sustained and superior validation accuracy for Scaf-GRPO throughout training.
RLVR with off-policy guidance.
To overcome the learning cliff, a phenomenon where a persistent lack of positive rewards renders difficult problems invisible to the learning gradient DAPO , researchers incorporate guidance from a “teacher” policy. The prevailing strategy is to provide the student model with a prefix of a “golden” trajectory and task it with generating the continuation yan2025learning ; Prefix-RFT ; StepHint ; BREAD ; ma2025learning . Different methods introduce variations on this theme. For instance, Luffy yan2025learning mixes a complete expert trajectory with multiple model-generated rollouts in one batch. Prefix-RFT Prefix-RFT employs a cosine decay schedule to adjust the length of the guiding prefix. More recently, StepHint StepHint provides multi-level hints of varying lengths, allowing the model to explore from multiple starting points. However, this prefix-continuation paradigm introduces challenges. It breaks policy consistency by mixing trajectories from two different distributions, which necessitates complex algorithmic patches yan2025learning ; Prefix-RFT ; StepHint . Furthermore, forcing the model down a predetermined path stifles exploration, limiting its ability to discover novel reasoning strategies. Our work provides effective guidance while circumventing these issues.
3  Methodology
Figure 3: Overview of the Scaf-GRPO framework. For a given query, the model generates multiple solutions. (Left) If any solution is correct, standard GRPO proceeds. (Right) If all solutions fail (the learning cliff), Scaf-GRPO initiates hierarchical hint-guided exploration. It injects progressively concrete in-prompt hints until a correct solution is found. This successful, minimally-guided trajectory replaces a failed one, restoring the learning gradient and enabling on-policy updates to resume.
Our framework, Scaffolded Group Relative Policy Optimization (Scaf-GRPO), illustrated in Figure 3, overcomes the learning cliff inherent in reinforcement learning by providing hierarchical, minimal, and progressive guidance. Unlike methods that alter the fundamental RL objective with off-policy data, Scaf-GRPO maintains the on-policy nature of GRPO. It operates by strategically augmenting the model’s rollout buffer when learning stagnates, ensuring that the learning signal is both meaningful and derived from the most efficient reasoning path the model can achieve with assistance. Our framework operates in two phases: an initial guidance exemption phase to diagnose “true-hard” problems, and a subsequent cyclical phase of hierarchical hint-guided exploration.
3.1  Preliminaries: Group Relative Policy Optimization (GRPO)
Group Relative Policy Optimization (GRPO) grpo is an on-policy RL algorithm for training LLMs that eliminates the need for a trainable value function. For a given prompt $q$, the policy $\pi_{\theta}$ generates a group of $N$ trajectories, $\mathcal{G}=\{o_{1},\dots,o_{N}\}$. After obtaining a terminal reward $R(o_{i})$ for each trajectory from a verifier, GRPO computes a normalized advantage $\hat{A}_{i}$ as: $\hat{A}_{i}=\frac{R(o_{i})-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}+\epsilon_{\text{std}}}$, where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in the group $\mathcal{G}$, and $\epsilon_{\text{std}}$ is a small constant for numerical stability. The policy is then updated by maximizing a clipped surrogate objective:
$J_{\text{GRPO}}(\theta)=\hat{\mathbb{E}}_{i,t}\left[\min\left(r_{i,t}(\theta)\hat{A}_{i},\,\text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right],$ (1)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|o_{i,<t},q)}{\pi_{\theta_{\text{old}}}(o_{i,t}|o_{i,<t},q)}$ is the probability ratio between the current and old policies, and $\epsilon$ is the clipping hyperparameter. The key limitation arises when all trajectories in $\mathcal{G}$ receive a zero reward, causing $\hat{A}_{i}$ to collapse to zero and stalling the learning process—the learning cliff.
3.2  The Scaf-GRPO Framework
Scaf-GRPO modifies the training process by strategically augmenting the trajectory group $\mathcal{G}$ when a learning cliff is detected. The process consists of a conditional batch construction procedure followed by the application of the standard GRPO loss.
Phase 1: Diagnosing true-hard problems.
A key principle of effective teaching is to avoid providing help when a learner can succeed independently. Not all initial failures indicate a fundamental capability gap; many are what we term pseudo-hard samples, arising from unfamiliarity with output formats or nascent reasoning skills. To address this, Scaf-GRPO incorporates a guidance exemption period, empirically set to the initial 15% of training steps. During this phase, the model attempts solutions purely through on-policy exploration. As shown in Figure 2, this period is characterized by a rapid decrease in zero-reward queries. We algorithmically determine when this independent learning has plateaued by monitoring the rate of solving zero-reward queries. Once this rate stagnates, any problem the model still consistently fails is classified as “true-hard,” making it a candidate for guidance. This ensures hints are reserved for genuine learning cliffs.
Phase 2: Hierarchical hint-guided exploration.
Once a problem is identified as “true-hard,” Scaf-GRPO activates its guidance mechanism using a pre-defined, three-tiered hint hierarchy, $H=\{H_{\text{knowledge}},H_{\text{planning}},H_{\text{solution}}\}$. The tiers offer distinct levels of guidance: (1) $H_{\text{knowledge}}$ (Knowledge Hint): Points to the key concept or formula required. (2) $H_{\text{planning}}$ (Planning Hint): Outlines a high-level strategic framework for the solution. (3) $H_{\text{solution}}$ (Solution Hint): Provides a concrete calculation step.
Figure 4: Prompt for hint injection.
To provide the minimal necessary guidance, the framework executes a deterministic search through this hierarchy, proceeding from the most abstract to the most concrete hint ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$). Within each tier, guidance is offered incrementally. The search terminates as soon as the model generates a correct solution, thereby identifying the minimal effective guidance required. A detailed description of this progressive exploration algorithm is provided in Appendix C.1.
On-policy batch augmentation and unified loss.
The core of Scaf-GRPO is its on-policy intervention, reactivating the learning signal during a learning cliff. When all initial trajectories $\mathcal{G}=\{o_{1},\dots,o_{N}\}$ from $\pi_{\theta}(\cdot|q)$ yield zero reward, the advantage $\hat{A}_{i}$ collapses, halting the gradient update. Scaf-GRPO intervenes by finding a minimal hint $h^{*}$ that enables policy $\pi_{\theta}$ to generate a successful trajectory $o_{h}^{*}\sim\pi_{\theta}(\cdot|q\oplus h^{*})$. This successful trajectory replaces a random failed trajectory $o_{j}\in\mathcal{G}$ to form an augmented group, $\mathcal{G}_{\text{final}}=(\mathcal{G}\setminus\{o_{j}\})\cup\{o_{h}^{*}\}$.
The key insight is that Scaf-GRPO does not alter the mathematical form of the GRPO loss function. Instead, it modifies the data used for the loss computation. The advantage calculation is performed on this conditionally augmented batch:
$\hat{A}_{i}^{\prime}=\frac{R(o_{i}^{\prime})-\mu_{\mathcal{G}_{\text{final}}}}{\sigma_{\mathcal{G}_{\text{final}}}+\epsilon_{\text{std}}}\quad\text{for }o_{i}^{\prime}\in\mathcal{G}_{\text{final}}.$ (2)
The learning objective remains the clipped surrogate objective, but it is now applied to the trajectories in $\mathcal{G}_{\text{final}}$. The probability ratio for a given trajectory $o_{i}^{\prime}\in\mathcal{G}_{\text{final}}$ at timestep $t$ is denoted as $r_{i,t}^{\prime}(\theta)$. The overall objective is:
$J_{\text{Scaf-GRPO}}(\theta)=\hat{\mathbb{E}}_{i,t}\left[\min\left(r_{i,t}^{\prime}(\theta)\hat{A}_{i}^{\prime},\text{clip}(r_{i,t}^{\prime}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}^{\prime}\right)\right],$ (3)
where the probability ratio $r_{i,t}^{\prime}(\theta)$ is critically computed with respect to the trajectory’s specific originating prompt:
$r_{i,t}^{\prime}(\theta)=\begin{cases}\frac{\pi_{\theta}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q)}{\pi_{\theta_{\text{old}}}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q)}&\text{if }o_{i}^{\prime}\in\mathcal{G}_{\text{final}}\text{ and }o_{i}^{\prime}\neq o_{h}^{*}\\ \frac{\pi_{\theta}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q\oplus h^{*})}{\pi_{\theta_{\text{old}}}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q\oplus h^{*})}&\text{if }o_{i}^{\prime}=o_{h}^{*}.\end{cases}$ (4)
This on-policy augmentation ensures the batch contains non-zero reward variance, restoring a meaningful advantage signal and allowing learning to resume on previously intractable problems.
Conservative nature and on-policy integrity.
A crucial property of Scaf-GRPO is its conservative nature; the framework does not alter the fundamental GRPO optimization objective but rather operates as a targeted data augmentation strategy. Its impact on the policy gradient can be formalized by analyzing two distinct cases based on the initial sampling results for a given prompt $q$.
In the first case, where at least one successful trajectory is generated initially ($\exists o_{i}\in\mathcal{G}$ such that $R(o_{i})>0$), the batch already contains a valid learning signal. The condition for intervention is not met, so the batch remains unchanged ($\mathcal{G}_{\text{final}}=\mathcal{G}$). Consequently, the objective function is mathematically identical to standard GRPO, ensuring our framework does not interfere when the model can learn on its own:
$J_{\text{Scaf-GRPO}}(\theta)\equiv J_{\text{GRPO}}(\theta).$ (5)
In the second case, the learning cliff scenario ($\forall o_{i}\in\mathcal{G},R(o_{i})=0$), standard GRPO fails. The uniform zero rewards cause the advantage calculation to collapse ($\mu_{\mathcal{G}}=0,\sigma_{\mathcal{G}}=0$), leading to a null advantage $\hat{A}_{i}=0$ and a vanishing policy gradient. Here, Scaf-GRPO intervenes by constructing the augmented batch $\mathcal{G}_{\text{final}}$. This restores the gradient by ensuring $\mu_{\mathcal{G}_{\text{final}}}>0$, which in turn guarantees a non-zero advantage signal $\hat{A}_{i}^{\prime}$. Critically, this intervention preserves the on-policy principle. Unlike off-policy methods that import trajectories from a different policy $\pi_{\phi}$ and require high-variance importance sampling corrections (e.g., using a ratio $\frac{\pi_{\theta}}{\pi_{\phi}}$), the guided trajectory $o_{h}^{*}$ is sampled directly from the current policy $\pi_{\theta}$. The probability ratio is therefore a standard on-policy ratio computed on a modified input, which is inherently more stable. This targeted, on-policy intervention transforms an unproductive, zero-gradient sample into a valuable learning opportunity without compromising the integrity of the optimization process.
Table 1: Overall performance on seven benchmarks. We compare our method, Scaf-GRPO, against vanilla GRPO baselines across diverse architectures, including the Qwen2.5 series, a non-Qwen model (Llama-3.2-8B-Instruct), and a specialized long-CoT model (DeepSeek-R1-Distill-Qwen-1.5B). Scores: pass@1 (%). Best results are in bold. The background color of Scaf-GRPO cells indicates performance change vs. Vanilla GRPO (green for improvement, red for decline).
Model AIME 24 AIME 25 AMC Minerva MATH-500 Olympiad Gaokao2023en Avg.
Qwen2.5-Math-1.5B
Qwen2.5-Math-1.5B 7.2 3.3 32.5 14.7 32.8 20.6 20.0 18.7
   Vanilla GRPO 13.3 10.0 47.5 28.3 72.2 34.8 57.4 37.6
   Scaf-GRPO 20.0 13.3 60.0 29.1 73.4 36.6 57.9 41.5
Qwen2.5-Math-7B
Qwen2.5-Math-7B 13.3 13.3 42.5 16.5 53.6 18.2 35.1 27.5
   Vanilla GRPO 30.0 13.3 60.0 33.4 75.8 41.3 62.6 45.2
   SimpleRL-Zero simplerlzoo 23.3 13.3 55.0 31.6 76.8 37.2 60.8 42.6
   Oat-Zero liu2025there 30.0 16.7 62.5 34.6 78.0 41.0 62.9 46.5
   LUFFY yan2025learning 33.3 16.7 62.5 33.8 75.2 41.7 62.7 46.6
   Scaf-GRPO 43.3 20.0 70.0 36.4 80.0 43.3 63.4 50.9
Qwen2.5-7B
Qwen2.5-7B 10.0 6.7 37.5 26.4 61.8 34.4 42.6 31.3
   Vanilla GRPO 10.0 10.0 50.0 38.5 77.6 40.4 64.2 41.5
   Scaf-GRPO 13.3 20.0 60.0 38.6 77.8 40.8 63.8 44.9
Llama-3.2-3B-Instruct
Llama-3.2-3B-Instruct 6.7 0.0 20.0 11.8 38.3 12.6 33.5 17.6
   Vanilla GRPO 13.3 0.0 35.0 18.7 51.8 18.3 45.7 26.1
   Scaf-GRPO 16.7 3.3 40.0 19.1 56.2 20.3 46.0 28.8
DeepSeek-R1-Distill-Qwen-1.5B
DeepSeek-R1-Distill-Qwen-1.5B 28.9 20.0 67.5 26.1 83.9 45.8 62.1 47.7
   Vanilla GRPO 30.0 21.1 67.5 30.1 83.9 50.2 71.4 50.6
   Scaf-GRPO 33.3 23.3 77.5 32.4 85.8 50.7 72.3 53.6
4  Experiments
4.1  Experimental Setups
Training dataset.
Our training data is derived from the DeepScaleR-Preview-Dataset deepscaler2025 . We employ a dynamic filtering strategy that aligns the dataset with each model’s initial capabilities. Based on preliminary evaluation, we classify problems as “Too Easy” (discarded), “Too Hard” (retained), or “Potentially Solvable” (50% subsampled). This curates a challenging yet tractable training set focused on the frontier of the model’s abilities. For this dataset, we generate our three-tiered hints by prompting the DeepSeek-R1 model DeepSeekR1 with ground-truth solution steps. Further details on our data filtering strategy and the hint generation process are provided in Appendix A.1 and Appendix A.2, respectively.
Models.
To demonstrate the general applicability of Scaf-GRPO, we conduct experiments across a diverse set of models, including: math-specialized models (Qwen2.5-Math-7B & 1.5B) to test in-domain performance; a general-purpose base model (Qwen2.5-7B) to assess skill acquisition; a different architecture (Llama-3.2-3B-Instruct) to verify model-agnosticism; and a Long-Chain-of-Thought model (DeepSeek-R1-Distill-Qwen-1.5B) to evaluate applicability to extended reasoning.
Baseline methods.
We benchmark Scaf-GRPO against three distinct classes of baselines: (1) Vanilla GRPO grpo , the standard on-policy algorithm without guidance. This serves as our baseline to quantify the gains from our scaffolding mechanism. (2) Leading GRPO implementations, including Simple-RL simplerlzoo and Oat-Zero liu2025there , to contextualize our performance against highly-optimized public benchmarks. (3) LUFFY yan2025learning , a representative of RLVR with off-policy guidance. This provides a direct comparison between the dominant prefix-continuation strategy and our in-prompt scaffolding approach.
Evaluation details.
We evaluate on diverse mathematics benchmarks, including GaoKao2023en chinesegaokao2024gaokao2023en , AIME24 aime24 , AIME25 aime25 , AMC amc23 , MATH-500 hendrycks2021math , and OlympiadBench he2024olympiadbench . To assess out-of-distribution (OOD) generalization, we evaluate on the scientific reasoning benchmark, GPQA-Diamond rein2024gpqa . For all benchmarks, we report pass@1 accuracy via greedy decoding. Vanilla GRPO is trained with our data and hyperparameters, and LUFFY on our data with its original parameters. For Simple-RL and Oat-Zero, we evaluate their publicly available weights.
Implementation details.
We train all models for 10 epochs using the verl framework sheng2025hybridflow , reporting results from the best-performing checkpoint. The maximum response length is 2048 tokens (8192 for the LongCoT model). Consistent with recent studies liu2025understandingr1zeroliketrainingcritical ; DAPO ; yan2025learning , we set the KL divergence penalty to zero to maximize policy exploration. A comprehensive list of hyperparameters is detailed in Appendix B.
Table 2: Ablation study on Scaf-GRPO’s key components using Qwen2.5-Math-7B model. The best performance is highlighted in bold. The “No Guidance” row serves as the vanilla GRPO baseline.
Hint Strategy AIME24 AIME25 AMC23 Minerva MATH-500 Olympiad Gaokao2023en Avg.
Scaf-GRPO (Full K →P →S) 43.3 20.0 70.0 36.4 80.0 43.3 63.4 50.9
w/o Progressive (Solution-Only) 40.0 13.3 65.0 36.2 78.6 43.7 62.3 48.4
w/o Knowledge Hint (P →S) 43.3 13.3 70.0 34.2 77.8 42.4 63.1 49.2
w/o Planning Hint (K →S) 43.3 16.7 62.5 35.0 79.4 40.0 63.6 48.6
w/o Solution Hint (K →P) 40.0 10.0 67.5 34.2 78.6 42.2 63.4 48.0
w/o Incremental Chunking 43.3 10.0 62.5 36.0 76.0 41.6 64.2 47.7
No Guidance (Vanilla GRPO) 30.0 13.3 60.0 33.4 75.8 41.3 62.6 45.2
4.2  Main Results
Comparison with GRPO.
As shown in Table 1, compared to the vanilla GRPO baseline, Scaf-GRPO achieves comprehensive and significant performance gains across all tested models. On the Qwen2.5-Math-7B model, Scaf-GRPO boosts the pass@1 score from 0.300 to 0.433 on AIME24, a relative improvement of 44.3%. These results provide strong evidence that our scaffolding mechanism effectively helps the model overcome the “learning cliff,” enabling it to tackle problems that were previously beyond its independent capabilities.
Comparison with other methods.
To contextualize Scaf-GRPO within the broader research landscape, we compare it against other leading methods in Table 1. Scaf-GRPO on Qwen2.5-Math-7B demonstrates a marked superiority, achieving an average score of 0.509. This performance represents a substantial improvement of 19.5% over Simple-RL and 9.5% over Oat-Zero. More importantly, Scaf-GRPO establishes a clear advantage over the prefix-continuation paradigm, outperforming LUFFY by 9.2%. This significant outperformance suggests that our in-prompt scaffolding strategy offers a more effective training alternative to prefix-continuation methods.
Generalization to non-Qwen architectures.
To verify that the benefits of Scaf-GRPO are not confined to a single model family, we extend our evaluation to the Llama-3.2-3B-Instruct model dubey2024llama . As detailed in Table 1, our framework demonstrates strong generalization. While vanilla GRPO provides a significant uplift over the base model, Scaf-GRPO achieves a further relative improvement of 10.3% in average performance. This confirms Scaf-GRPO is a model-agnostic method, capable of enhancing reasoning abilities beyond the Qwen series.
Applicability to LongCoT models.
We further investigate the efficacy of Scaf-GRPO on models optimized for Long Chain-of-Thought (LongCoT) reasoning, using the specialized DeepSeek-R1-Distill-Qwen-1.5B model. The results in Table 1 show that Scaf-GRPO effectively enhances this already capable baseline, delivering a 5.9% relative performance gain over vanilla GRPO. This demonstrates our framework’s versatility in scaffolding not only standard-length solutions but also the extensive derivations characteristic of LongCoT models.
4.3  Ablation Study
We conduct a series of ablation studies on the Qwen2.5-Math-7B model (see Table 2).
Necessity of the guidance exemption period.
To validate our guidance exemption period (Phase 1), we tested a variant applying scaffolding from the start of training. As detailed in Appendix F, this resulted in a 9.2% relative performance drop for the Qwen2.5-Math-7B model compared to our full framework. This confirms that an initial autonomous learning period is crucial to prevent over-reliance on hints and foster more robust, independent reasoning.
Efficacy of progressive & hierarchical guidance.
Our methodology is founded on the hypothesis that progressive guidance, from abstract concepts to concrete steps, is superior to simply providing a direct solution. To test this, we evaluate a “Solution-Only” variant that bypasses the hierarchy and immediately provides the most concrete hint. This results in a significant performance degradation of 4.9% compared to the full model. This confirms our hypothesis: compelling the model to first engage with higher-level reasoning fosters more generalizable skills.
Figure 5: Evolution of reasoning from guidance to autonomy. The model progresses from imitating a concrete hint (a) to applying abstract knowledge (b), ultimately achieving (c) autonomous problem-solving by internalizing the guided skills.
Justifying the completeness of the hint hierarchy.
We design a three-tiered hint structure (K→P→S) assuming each layer serves a unique function. To verify this, we systematically removed one layer at a time. As shown in Table 2, every removal degrades performance. The most severe degradation, a 5.7% drop, occurs when the final “Solution” hint is removed (the K→P variant). This highlights the dual role of the hierarchy: abstract hints encourage high-level reasoning, while concrete hints serve as an essential fallback. The superior performance of the full K→P→S model demonstrates that the layers are complementary, not redundant.
Efficacy of incremental guidance.
A core principle of Scaf-GRPO is to provide the minimal necessary support by delivering hints incrementally. We test this against a “Full Hint” variant, which provides the entire content of a hint level at once. This non-incremental approach collapses performance by 6.3% compared to the incremental one. This decline validates our strategy: minimal, incremental intervention is critical for preserving model autonomy and preventing over-reliance.
4.4  Further Analysis
Confronting the learning cliff.
Figure 2 visualizes Scaf-GRPO’s advantage. In Figure 2(a), we plot the number of “zero-reward” problems per batch. The count for both methods drops sharply at the start of training. However, the vanilla GRPO curve quickly flattens, defining the learning cliff: a point where the baseline can no longer extract a learning signal from a persistent set of “true-hard” problems. In contrast, Scaf-GRPO’s scaffolding activates, enabling the model to consistently learn from these problems and continue reducing the zero-reward count. This directly impacts validation performance (Figure 2(b)). By turning intractable problems into learning opportunities, Scaf-GRPO achieves a higher, steadily improving validation score while the baseline stagnates.
Table 3: Impact of data filtering on Scaf-GRPO vs. Vanilla GRPO. Both methods were trained on the full dataset (Original) and a harder subset (Filtered). The best performance is highlighted in bold. Scores are pass@1 (%).
Data Method AIME24 AIME25 AMC23 MATH-500 Olympiad Avg.
Qwen2.5-Math-1.5B
Original Vanilla GRPO 13.3 6.7 52.5 68.6 31.4 34.5
Original Scaf-GRPO 20.0 10.0 55.0 73.2 36.4 38.9
Filtered Vanilla GRPO 13.3 10.0 47.5 72.2 34.8 35.6
Filtered Scaf-GRPO 20.0 13.3 60.0 73.4 36.6 40.7
Qwen2.5-Math-7B
Original Vanilla GRPO 30.0 16.7 60.0 74.4 38.5 43.9
Original Scaf-GRPO 33.3 16.7 70.0 79.0 43.0 48.4
Filtered Vanilla GRPO 30.0 13.3 60.0 75.8 41.3 44.1
Filtered Scaf-GRPO 43.3 20.0 70.0 80.0 43.3 51.3
Internalizing skills beyond imitation.
Scaf-GRPO succeeds by fostering skill acquisition, not just imitation. Figure 5 illustrates this on a true-hard problem. The model’s reasoning evolves from imitating concrete “Solution Hints” to following strategic “Planning Hints” and finally to applying abstract “Knowledge Hints.” The key evidence for true learning appears when the model later revisits the problem. Without any hints, it successfully solves it by combining skills from its prior guided experiences. This demonstrates that hints serve as a tool to build lasting abilities, not just to achieve temporary success. This process of skill-building on hard problems overcomes the learning cliff.
Aligning data difficulty with model capacity.
To validate our data filtering strategy, we train models on both the complete dataset and our filtered subset. As detailed in Table 3, the harder, filtered data yields a marginal 0.5% relative gain for vanilla GRPO but a substantial 6.0% boost for Scaf-GRPO on Qwen2.5-Math-7B. This disparity underscores that exposing a model to difficult problems is insufficient. A challenging curriculum is most effective when paired with a robust learning framework like Scaf-GRPO, which converts these challenges into learning opportunities.
Generalization to out-of-distribution tasks.
To verify that Scaf-GRPO cultivates robust reasoning skills rather than in-domain pattern matching, we evaluate its generalization on the out-of-distribution GPQA-Diamond benchmark, which features expert-level scientific questions. Scaf-GRPO demonstrates strong OOD performance, elevating the Qwen2.5-Math-7B model’s pass@1 score to 37.3%. This represents a significant improvement over the base model (24.7%) and the vanilla GRPO baseline (32.3%), and matches the performance of the strong prefix-based LUFFY baseline. This result indicates that the problem-solving abilities fostered by our scaffolding approach are fundamental and transfer effectively to novel domains, a key attribute for building more general AI reasoners.
Table 4: OOD performance (pass@1,%) of Qwen2.5-Math-7B on the GPQA-Diamond benchmark.
Model GPQA-Diamond
Qwen2.5-Math-7B
Base Model 24.7
Vanilla GRPO 32.3
SimpleRL-Zero simplerlzoo 33.3
Oat-Zero liu2025there 33.3
LUFFY yan2025learning 37.3
Scaf-GRPO (Ours) 37.3
5  Discussion and Conclusion
Limitations.
The practical deployment of Scaf-GRPO is subject to two main considerations. First, its efficacy currently relies on the availability of a high-quality, tiered hint hierarchy. Generating these structured hints requires a non-trivial data preparation effort. Second, the framework is principally designed for tasks with verifiable solutions and structured reasoning paths, such as mathematics. Its applicability to more open-ended, subjective domains like creative writing is less direct.
Future work.
Future research could focus on automating hint generation to enhance the framework’s scalability. We also plan to explore adaptive scaffolding mechanisms where guidance dynamically adjusts to the model’s improving proficiency, thereby personalizing the learning process.
Conclusion.
In this work, we introduce Scaf-GRPO, a training framework that overcomes the “learning cliff” in reinforcement learning for large language models. By providing hierarchical hints in the prompt, Scaf-GRPO offers scaffolding for models to solve problems beyond their reach. This on-policy guidance preserves exploratory autonomy and mitigates the distributional consistency issues inherent in prefix-continuation methods. Our experiments show Scaf-GRPO significantly outperforms vanilla GRPO and strong prefix-based baselines across challenging mathematics benchmarks. This framework enables models to learn from previously intractable problems, establishing a more effective path toward autonomous reasoning.
Appendix A Dataset and Benchmark Details
A.1  Training Data Source and Filtering
Our training data is derived from the DeepScaleR-Preview-Dataset [21], a comprehensive collection of 40k mathematical problems. Its contents are sourced from AIME, AMC, MATH [26], Still [6], and Omni-MATH [31]. To maximize training efficiency and target problems most conducive to learning, we implement a dynamic data filtering strategy tailored to each model’s capabilities. This strategy categorizes problems based on the model’s initial performance, assessed through 8 independent sampling attempts for each problem. These samples are generated using nucleus sampling with a temperature of 1.0 (top-p=1.0, top-k=-1) and a maximum length of 2048 tokens for both the prompt and the response. For our LongCoT model, this response token was increased to 8192 tokens to accommodate its generation style. Based on the outcomes, problems are categorized as follows:
  • Too Easy: Problems solved correctly in all 8 attempts are excluded, as they offer minimal learning value.
  • Potentially Solvable: Problems solved in 1 to 7 of the attempts are considered to be within the model’s learning-rich sweet spot. We randomly sample 50% of these for inclusion.
  • Too Hard: Problems that fail in all 8 attempts are retained, as they are the primary candidates for our scaffolding mechanism.
This filtering process results in a final training dataset where approximately 50% of the problems are from the “Potentially Solvable” category and the remaining 50% are "Too Hard". This curated dataset is challenging yet tractable, maximizing the efficiency of the training process for both the baseline GRPO and our Scaf-GRPO framework.
A.2  Hint Generation for Training Data
The hierarchical hints ($H_{\text{knowledge}},H_{\text{planning}},H_{\text{solution}}$) are the cornerstone of Scaf-GRPO’s guidance mechanism. To create them, we perform a one-time, offline preprocessing step using the powerful DeepSeek-R1 model [3]. For each problem in our curated training set, we provide the model with both the problem statement and its ground-truth solution trace.
We then employ a highly structured prompt, detailed in Appendix D, which is engineered not only to decompose the solution into our three-tiered hierarchy but also to enforce a crucial internal structure. Specifically, the prompt compels the model to generate exactly four numbered, progressive items for each category. These items are designed to build upon one another, creating four distinct levels of guidance. For instance, the four items for an $H_{\text{planning}}$ hint might represent: (1) the first step of the plan, (2) the second step, (3) the third step, and (4) the fourth step. This structured, multi-level design within each hint category enables the fine-grained, progressive exploration central to our method. The entire process ensures a consistent and high-quality set of structured hints, and the scripts will be included in our code release.
A.3  Evaluation Benchmarks
We evaluate all models on a diverse suite of seven challenging mathematics benchmarks and the GPQA-Diamond benchmark to ensure a robust and comprehensive assessment of their reasoning abilities. Table 5 provides details for each benchmark used. The mathematics datasets span various difficulty levels and mathematical domains, from high-school competition problems to Olympiad-level challenges, providing a rigorous testbed for advanced reasoning. GPQA-Diamond serves as a crucial Out-of-Distribution (OOD) benchmark, testing generalization to expert-level scientific questions outside the training domain. All benchmarks are publicly available.
A.4  Motivation for Dataset Selection
Our choice of the DeepScaleR-Preview-Dataset for training is deliberate. This decision is motivated by two key factors. First, the dataset’s broad scope, encompassing problems of varying difficulty, provides the rich and diverse material necessary for our filtering strategy to be effective. Second, its successful application in prior research [6] establishes it as a robust and relevant foundation for training advanced reasoning models.
For evaluation, our selection of benchmarks is designed for a rigorous and multifaceted assessment. The suite primarily consists of challenging competition-level benchmarks (AIME24/25, AMC, MATH, OlympiadBench, Minerva) and a standardized national exam (GaoKao2023), covering a broad spectrum of mathematical reasoning. Crucially, to measure out-of-distribution (OOD) generalization, we include the GPQA-Diamond benchmark [28]. As GPQA consists of graduate-level questions whose style and domain are distinct from our training data, strong performance on this benchmark indicates that Scaf-GRPO fosters genuine reasoning skills rather than mere pattern memorization of the training distribution.
Benchmark Description Citation # Problems
AIME24 American Invitational Mathematics Examination 2024 [23] 30
AIME25 American Invitational Mathematics Examination 2025 [24] 30
AMC American Mathematics Competitions 2023 [25] 25
MATH-500 A subset from the MATH test set [26] 500
GaoKao2023en Chinese National College Entrance Exam 2023 [22] 385
OlympiadBench Math Olympiad-level problems [27] 675
Minerva A specialized dataset to evaluate quantitative and scientific reasoning abilities [32] 272
GPQA-Diamond Expert-level questions across biology, physics, and chemistry [28] 198
Table 5: Details of evaluation benchmarks used in our experiments.
Appendix B Experimental Setup Details
B.1  Computing Infrastructure
All experiments are conducted on a high-performance computing cluster. The specific hardware and software configurations are as follows:
  • Hardware: All models are trained and evaluated on servers equipped with 8 NVIDIA A100 (80GB) GPUs.
  • Software: The operating system is Ubuntu 22.04. Key software libraries and their versions include PyTorch 2.6.0, Transformers 4.51.1, and CUDA 12.4.
  • Framework: Our implementation is built upon the verl (0.4.1) framework [29], a robust and efficient library designed for large-scale reinforcement learning with LLMs.
B.2  Hyperparameter Details
Our experimental setup is carefully configured for performance and reproducibility. The final hyperparameter configuration is detailed comprehensively in Table 6. These settings were applied consistently across all experiments to ensure a fair comparison. However, to accommodate models specialized in Long Chain-of-Thought (LongCoT) reasoning, we increased the maximum response length to 8192 tokens for those specific tasks.
Hyperparameter Value
Optimization & Training
Learning Rate (LR) $1\times 10^{-6}$
Optimizer AdamW
Weight Decay 0.0
Batching Strategy
Global Batch Size 256
PPO Mini-batch Size 64
Micro-batch Size (per GPU) 16
Validation Batch Size 512
RL Algorithm (GRPO)
Rollouts per Query ($N$) 8
GRPO Clip Epsilon ($\epsilon$) 0.2
KL Divergence Penalty ($\beta$) 0.0
Entropy Coefficient 0.0
Generation & Tokenization
Rollout Temperature 1.0
Max Response Tokens 2048
Infrastructure & Scheduling
Nodes 1
GPUs per Node 8
VLLM GPU Memory Utilization 0.8
Table 6: Comprehensive list of key hyperparameters for training and generation.
B.3  Implementation of Baseline Methods
To ensure a rigorous and fair comparison, we carefully implement or utilize baselines as follows:
Vanilla GRPO.
This is our primary control. We train a vanilla GRPO model using our verl framework. It is configured with the exact same filtered dataset and hyperparameters as Scaf-GRPO, allowing us to cleanly isolate the performance contribution of our scaffolding mechanism.
LUFFY.
To compare against the dominant prefix-continuation paradigm, we train LUFFY [2] using its official public implementation and its original, recommended hyperparameter settings. To ensure a fair comparison of methodological effectiveness, we train it on our high-quality filtered dataset. This directly contrasts their off-policy guidance with our in-prompt scaffolding on identical data.
Simple-RL and Oat-Zero.
For other leading methods like Simple-RL [7] and Oat-Zero [20], we do not perform any implementation or retraining. Instead, we evaluate their officially released, publicly available model checkpoints directly. All reported results for these models are obtained by running them through our unified evaluation pipeline, ensuring a consistent and fair comparison against established state-of-the-art work.
B.4  Evaluation Metrics Details
Our primary evaluation metric is pass@1, which measures the percentage of problems for which a model generates a correct solution in a single attempt. This metric is chosen for its straightforwardness and its status as a standard for evaluating definitive problem-solving capabilities. For all evaluations, we use greedy decoding to generate one complete solution trace for each problem.
The verification process is tailored to the benchmark type to ensure maximum rigor and fairness.
  • For all mathematical reasoning benchmarks, we employ the “symeval” library [33], specifically its EvaluatorMathBatch module, to determine correctness. This approach moves beyond simple string comparison by using a sophisticated pipeline that combines regular expressions for robust answer extraction with SymPy for symbolic mathematical evaluation. This allows for the accurate verification of complex answers, including matrices, intervals, and symbolic expressions.
  • For the out-of-distribution GPQA-Diamond benchmark, we utilize the EleutherAI’s lm-evaluation-harness [34] to ensure a fair and standardized assessment. This widely adopted framework provides a consistent testing environment for generative models. We use its implementation of the gpqa-diamond task to compute the pass@1 score, thereby maintaining metric consistency while leveraging a community-standardized evaluation harness for OOD generalization.
Appendix C Methodology Details
This section provides a granular description of the Progressive Exploration and Replacement Algorithm, which is central to how Scaf-GRPO overcomes the learning cliff by strategically providing minimal guidance during training.
C.1  The Progressive Exploration and Replacement Algorithm
When a “true-hard” problem triggers the hierarchical hint-guided exploration phase, Scaf-GRPO executes a deterministic, multi-level search algorithm to find the minimal effective hint. The algorithm’s goal is to provide just enough information for the model to succeed, thereby maximizing its independent reasoning.
The algorithm leverages the pre-generated, four-level progressive hint structure, detailed in Appendix A.2. It systematically searches through the hint categories in order of decreasing abstraction ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$) and, within each category, through the four levels of increasing detail.
Let $h_{c}^{i}$ denote the $i$-th hint item for a category $c\in\{\text{knowledge, planning, solution}\}$. The cumulative hint provided to the model at level $l\in\{1,2,3,4\}$ is the union of the first $l$ items, denoted as $C_{c}^{l}=\bigcup_{i=1}^{l}\{h_{c}^{i}\}$. The search process for a single problem is as follows:
  1. (1)
    Iterate through Hint Categories: For each category $c$ in the sequence (“knowledge”, “planning”, “solution”):
    1. (a)
      Iterate through Hint Levels: For each level $l$ from 1 to 4:
      1. i.
        Construct an augmented prompt by injecting the cumulative hint $C_{c}^{l}$.
      2. ii.
        Generate a new solution on-policy using this augmented prompt.
      3. iii.
        If the generated solution is correct, the search successfully terminates. The trajectory produced with hint $C_{c}^{l}$ replaces one of the failed trajectories in the batch. The algorithm then concludes for this problem.
  2. (2)
    Handle Intractable Case: If the nested loops complete without finding a correct solution (i.e., even the most detailed hint $C_{\text{solution}}^{4}$ fails), the problem is deemed intractable for the current training step. No replacement occurs, and the algorithm concludes for this problem, leaving the original all-failure group in the batch.
This structured and exhaustive search ensures that if a solution is reachable with any level of guidance, the framework will find it using the most abstract and minimal hint possible, thereby preserving the on-policy learning signal for “true-hard” problems.
Appendix D Prompt Design in Scaf-GRPO
The effectiveness of Scaf-GRPO relies on two distinct but complementary types of structured prompts: those for generating the hierarchical hints, and those for injecting these hints during training.
D.1  Hint Generation Prompt
To systematically create our tiered hints, we provide a powerful teacher model (DeepSeek-R1) with a structured prompt. For each problem-solution pair in our dataset, this prompt instructs the teacher model to decompose the solution into our three-tiered hierarchy ($H_{\text{knowledge}},H_{\text{planning}},H_{\text{solution}}$). This semi-automated process is a critical preprocessing step that ensures a consistent and high-quality hint dataset. The exact prompt template used is below.
Prompt for hint injection **[ROLE & GOAL]** {internallinenumbers*} You are an expert AI assistant specializing in problem-solving methodology and knowledge engineering. Your task is to analyze a given problem and its ground-truth solution, and then generate a structured breakdown of the reasoning process with a high degree of granularity. **[INPUT]** I will provide you with a "Problem" and its "Ground-Truth Solution". **[INSTRUCTIONS]** {internallinenumbers*} Based on the provided input, you must generate exactly THREE components. For each component, you MUST generate **a minimum of 4 numbered items**. {internallinenumbers*} - If a natural breakdown results in fewer than 4 items, you must **subdivide the existing steps into more detailed, finer-grained sub-steps** to meet the requirement. For example, a single calculation step can be broken down into ’substituting values’, ’performing the operation’, and ’stating the result’. {internallinenumbers*} 1. **Planning Skeleton**: Extract a high-level planning skeleton. This should be a concise, ordered list of the key reasoning steps and the overall strategy used to reach the solution. Do not include detailed calculations, just the logical flow. Break it down into at least 4 detailed steps. {internallinenumbers*} 2. **Knowledge Components**: Identify at least 4 essential knowledge components (like facts, definitions, theorems, lemmas, or formulas) required to solve the problem. List each component clearly in a numbered list. {internallinenumbers*} 3. **Solution Breakdown**: Divide the original Ground-Truth Solution into a numbered list of semantically coherent steps or chunks.There should be at least 4 steps or chunks. Each item in the list should be a direct quote or a faithful summary of a part of the original solution text. **[OUTPUT FORMAT]** {internallinenumbers*} You MUST provide your response in the following structured format. Ensure each section contains at least 4 items. <PLANNING_SKELETON> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </SOLUTION_BREAKDOWN> **[EXAMPLE]** --- BEGIN EXAMPLE --- # ... (Example Problem and Solution are the same) --- END EXAMPLE --- **[EXPECTED OUTPUT FOR THE EXAMPLE]** (This example now demonstrates the required granularity with >= 4 items per section) <PLANNING_SKELETON> 1. Identify the geometric shape (right-angled triangle) and the goal (find the hypotenuse). 2. Recall the relevant mathematical theorem that connects the sides of a right-angled triangle. 3. Formulate the equation by substituting the given side lengths (base and height) into the theorem. 4. Execute the arithmetic calculation to find the square of the hypotenuse. 5. Perform the final step of taking the square root to isolate the length of the hypotenuse. </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. Theorem: Pythagorean Theorem (a² + b² = c² for a right-angled triangle). 2. Definition: Hypotenuse (The longest side of a right-angled triangle, opposite the right angle). 3. Concept: Right-angled Triangle (A triangle with one angle measuring 90 degrees). 4. Mathematical Operation: Square Root (The inverse operation of squaring a number). </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> {internallinenumbers*} 1. To find the hypotenuse of a right-angled triangle, we can use the Pythagorean theorem, which states that a² + b² = c², where a and b are the lengths of the two shorter sides (legs) and c is the length of the hypotenuse. 2. The given values are a = 4 cm and b = 3 cm. 3. Substituting these into the formula gives: 4² + 3² = 16 + 9 = 25. 4. This means c² = 25. Taking the square root of both sides results in c = 5 cm. </SOLUTION_BREAKDOWN> --- **[TASK]** Now, please process the following problem and solution, strictly following all instructions. **Problem**: {question} **Ground-Truth Solution**: {solution}
D.2  Inject Hints Prompt
During the Hierarchical Hint-Guided Exploration phase, when the model fails to solve a “true-hard” problem, Scaf-GRPO injects a hint directly into the input prompt. This approach is fundamental to our on-policy methodology, as it reframes the problem for the model rather than forcing it to continue a partial, off-policy trajectory. The design of this prompt is crucial: it explicitly informs the model that it is receiving guidance, ensuring that the model processes both the problem and the hint under a single, unified policy, thereby avoiding the distributional shifts common in prefix-continuation methods.
The exact prompt template used for hint injection is shown below.
[…]
Appendix E Formal Algorithmic and Mathematical Description of Scaf-GRPO
This section provides a formal mathematical and algorithmic description of the Scaffolded Group Relative Policy Optimization (Scaf-GRPO) framework. We formalize the two-phase training process and detail the construction of the loss function, particularly for the hierarchical hint-guided exploration phase.
E.1  Preliminaries: The Standard GRPO Objective
We begin by restating the core Group Relative Policy Optimization (GRPO) objective. For a given prompt $q$, the policy $\pi_{\theta}$ generates a group of $N$ trajectories, $\mathcal{G}=\{o_{1},\dots,o_{N}\}$, where each trajectory is sampled from the policy, $o_{i}\sim\pi_{\theta}(\cdot|q)$. Each trajectory receives a terminal reward $R(o_{i})$ from an external verifier.
The normalized advantage for each trajectory $o_{i}$ in the group $\mathcal{G}$ is calculated as:
$\hat{A}_{i}=\frac{R(o_{i})-\mu_{\mathcal{G}}}{\sigma_{\mathcal{G}}+\epsilon_{\text{std}}}$ (6)
where $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ are the mean and standard deviation of rewards in $\mathcal{G}$, and $\epsilon_{\text{std}}$ is a small constant for numerical stability.
The standard GRPO objective, which is maximized during training, is a clipped surrogate objective defined as the empirical expectation over trajectories and timesteps:
$J_{\text{GRPO}}(\theta)=\hat{\mathbb{E}}_{i,t}\left[\min\left(r_{i,t}(\theta)\hat{A}_{i},\,\text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\right)\right]$ (7)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|o_{i,<t},q)}{\pi_{\theta_{\text{old}}}(o_{i,t}|o_{i,<t},q)}$ is the probability ratio for the token at timestep $t$ of trajectory $i$ between the current and old policies, and $\epsilon$ is the clipping hyperparameter. The “learning cliff” phenomenon, a key challenge addressed by our work, occurs when $R(o_{i})=0$ for all $i\in\{1,\dots,N\}$. In this scenario, $\mu_{\mathcal{G}}$ and $\sigma_{\mathcal{G}}$ become zero, causing the advantage $\hat{A}_{i}$ to collapse to zero for the entire group and stall the learning process.
E.2  The Scaf-GRPO Training Process: A Two-Phase Formulation
Scaf-GRPO modifies the training process by strategically augmenting the trajectory group $\mathcal{G}$ when a learning cliff is detected. The process consists of a conditional batch construction procedure followed by the application of the standard GRPO loss.
Let $t$ denote the current training step and $T_{\text{exempt}}$ be the step at which the guidance exemption period ends. The core logic is detailed in Figure 6.
Figure 6: Overview of the Scaffolded Group Relative Policy Optimization (Scaf-GRPO) Algorithm.
The function $\text{SearchHierarchicalHints}(q,\pi_{\theta})$ represents the deterministic, multi-level search described in Section 3.3 and Appendix D. It iterates through the pre-defined hint hierarchy $\mathcal{H}=\{\mathcal{H}_{\text{knowledge}},\mathcal{H}_{\text{planning}},\mathcal{H}_{\text{solution}}\}$ to find the first hint $h^{*}$ that enables the policy $\pi_{\theta}$ to generate a successful trajectory $o_{h}^{*}\sim\pi_{\theta}(\cdot|q\oplus h^{*})$, where $\oplus$ denotes the concatenation of the hint into the prompt. If no hint leads to a solution, it returns $(\text{null},\text{null})$.
E.3  The Unified Scaf-GRPO Loss Function
The core insight of Scaf-GRPO is that it does not alter the mathematical form of the GRPO loss function. Instead, it modifies the data distribution used for the loss computation by conditionally augmenting the batch.
Let $\mathcal{G}_{\text{final}}$ denote the group of trajectories returned by Figure 6 for a given prompt $q$ at step $t$. This group is composed of trajectories sampled under one of two conditions:
  1. (1)
    Standard Generation: All $N$ trajectories are from $\pi_{\theta}(\cdot|q)$. This occurs if the learning cliff is not triggered or if the training is within the exemption period.
  2. (2)
    Augmented Generation: N1N-1 trajectories are from $\pi_{\theta}(\cdot|q)$ (with zero reward), and one trajectory, $o_{h}^{*}$, is from $\pi_{\theta}(\cdot|q\oplus h^{*})$ (with positive reward). This occurs only when the learning cliff is triggered post-exemption and the hint search is successful.
The Scaf-GRPO loss function is therefore defined by applying the standard GRPO objective to this conditionally constructed batch. First, the advantage is computed on the final group $\mathcal{G}_{\text{final}}$:
$\hat{A}_{i}^{\prime}=\frac{R(o_{i}^{\prime})-\mu_{\mathcal{G}_{\text{final}}}}{\sigma_{\mathcal{G}_{\text{final}}}+\epsilon_{\text{std}}}\quad\text{for }o_{i}^{\prime}\in\mathcal{G}_{\text{final}}$ (8)
The overall objective is then:
$J_{\text{Scaf-GRPO}}(\theta)=\hat{\mathbb{E}}_{i,t}\left[\min\left(r_{i,t}^{\prime}(\theta)\hat{A}_{i}^{\prime},\,\text{clip}(r_{i,t}^{\prime}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}^{\prime}\right)\right]$ (9)
where the probability ratio $r_{i,t}^{\prime}(\theta)$ for each trajectory $o_{i}^{\prime}\in\mathcal{G}_{\text{final}}$ is critically computed with respect to its specific originating prompt:
$r_{i,t}^{\prime}(\theta)=\begin{cases}\frac{\pi_{\theta}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q)}{\pi_{\theta_{\text{old}}}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q)}&\text{if }o_{i}^{\prime}\in\mathcal{G}_{\text{final}}\text{ and }o_{i}^{\prime}\neq o_{h}^{*}\\ \frac{\pi_{\theta}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q\oplus h^{*})}{\pi_{\theta_{\text{old}}}(o_{i,t}^{\prime}|o_{i,<t}^{\prime},q\oplus h^{*})}&\text{if }o_{i}^{\prime}=o_{h}^{*}.\end{cases}$ (10)
By reformulating the batch rather than the loss, Scaf-GRPO ensures that when a learning signal is absent ($\hat{A}_{i}=0$), a new signal is injected by providing a single successful, minimally-guided trajectory. This intervention re-establishes a non-zero reward variance within the group, reactivates the advantage calculation, and enables learning to resume on previously intractable problems, thereby directly overcoming the learning cliff.
E.4  Conservative Nature and Preservation of the On-Policy Objective
A crucial property of Scaf-GRPO is that it does not alter the fundamental GRPO optimization objective. Instead, it operates as a conservative data augmentation strategy that activates only under the specific condition of a learning cliff. We can formalize the framework’s impact on the policy gradient by analyzing two distinct cases based on the initial on-policy sampling results for a given prompt $q$.
Case 1: At least one successful trajectory ($\exists o_{i}\in\mathcal{G}$ such that $R(o_{i})>0$).
In this scenario, the initial group of trajectories $\mathcal{G}$ already contains a non-uniform reward signal, meaning $\mu_{\mathcal{G}}>0$ and $\sigma_{\mathcal{G}}\geq 0$. The condition for triggering the hierarchical hint-guided exploration is not met. Consequently, the final batch used for the update is the original batch, $\mathcal{G}_{\text{final}}=\mathcal{G}$. The Scaf-GRPO objective function is therefore mathematically identical to the standard GRPO objective:
$J_{\text{Scaf-GRPO}}(\theta)\equiv J_{\text{GRPO}}(\theta)$ (11)
In the most frequent training scenarios where the model has some capacity to solve the problem, our framework makes no modifications and is equivalent to vanilla GRPO.
Case 2: All trajectories fail ($\forall o_{i}\in\mathcal{G}$, $R(o_{i})=0$).
This is the learning cliff scenario. In standard GRPO, the rewards are uniform and zero, causing the advantage calculation to collapse: $\mu_{\mathcal{G}}=0$ and $\sigma_{\mathcal{G}}=0$, leading to $\hat{A}_{i}=0$ for all trajectories. The resulting policy gradient for this prompt is zero, and no learning occurs.
Scaf-GRPO intervenes by constructing an augmented batch $\mathcal{G}_{\text{final}}$. This batch consists of N1N-1 of the original failed trajectories and one new, successful trajectory $o_{h}^{*}\sim\pi_{\theta}(\cdot|q\oplus h^{*})$. The crucial insight is that this new trajectory is generated on-policy by the current policy $\pi_{\theta}$, conditioned on the hint-augmented prompt.
The key benefits of this intervention are:
  1. (1)
    Gradient Restoration. The augmented batch $\mathcal{G}_{\text{final}}$ now contains at least one trajectory with a positive reward. This ensures that $\mu_{\mathcal{G}_{\text{final}}}>0$ and $\sigma_{\mathcal{G}_{\text{final}}}>0$, which in turn guarantees a non-zero advantage signal $\hat{A}_{i}^{\prime}$ for the trajectories in the group. Learning is effectively restored where it would have stalled.
  2. (2)
    Preservation of the On-Policy Principle. Unlike off-policy methods that mix trajectories from a different policy $\pi_{\phi}$ and require importance sampling corrections (e.g., $\frac{\pi_{\theta}}{\pi_{\phi}}$) to account for the distributional shift, Scaf-GRPO’s guided trajectory $o_{h}^{*}$ is sampled directly from the current policy $\pi_{\theta}$. Therefore, the probability ratio $r_{i,t}^{\prime}(\theta)$ is a standard on-policy ratio computed at each timestep. This avoids the high variance and potential instability associated with off-policy corrections, ensuring that the learning signal remains stable and directly attributable to the current policy’s capabilities.
In summary, Scaf-GRPO does not introduce any harmful bias or modification to the GRPO objective. It is a targeted intervention that is inactive when a valid learning signal already exists. When the learning signal vanishes, it provides a constructive, on-policy gradient by minimally augmenting the task, thereby transforming an unproductive training sample into a valuable learning opportunity without compromising the integrity of the on-policy optimization process.
Appendix F Ablation on the Guidance Exemption Period
[…]
Applying scaffolding prematurely to these pseudo-hard cases is suboptimal. It fosters dependency, teaching the model to rely on external hints for problems it could have solved through autonomous exploration. This intervention prevents the model from correcting its own foundational errors and developing robust, independent problem-solving habits. Our hypothesis is that Phase 1 allows the model to overcome these basic execution issues on its own, ensuring that the scaffolding mechanism is reserved for addressing genuine reasoning challenges.
To test this hypothesis, we compare our full Scaf-GRPO framework against a variant where scaffolding is activated from the very beginning of training (“Scaf-GRPO w/o Phase 1”). The performance of both methods, alongside the vanilla GRPO baseline, is presented in Table 7 for the Qwen2.5-Math models.
Table 7: Ablation study on the necessity of the guidance exemption period (Phase 1). We compare the full Scaf-GRPO framework against vanilla GRPO and a Scaf-GRPO variant without the initial exemption phase. Scores: pass@1 (%). Best results are in bold.
Model AIME 24 AIME 25 AMC Minerva MATH-500 Olympiad Gaokao2023en Avg.
Qwen2.5-Math-1.5B
Vanilla GRPO 13.3 10.0 47.5 28.3 72.2 34.8 57.4 37.6
Scaf-GRPO (w/o Phase 1) 10.0 10.0 57.5 27.5 71.4 36.3 57.9 38.7
Scaf-GRPO 20.0 13.3 60.0 29.1 73.4 36.6 57.9 41.5
Qwen2.5-Math-7B
Vanilla GRPO 30.0 13.3 60.0 33.4 75.8 41.3 62.6 45.2
Scaf-GRPO (w/o Phase 1) 23.3 13.3 70.0 34.2 78.4 41.2 63.1 46.2
Scaf-GRPO 43.3 20.0 70.0 36.4 80.0 43.3 63.4 50.9
The results in Table 7 confirm our hypothesis. While activating scaffolding from the start yields a significant improvement over the vanilla GRPO baseline, its performance is notably inferior to the complete Scaf-GRPO framework. This performance gap underscores the criticality of the guidance exemption period. By allowing the model an initial phase of unguided learning, we prevent over-reliance on hints and ensure that our scaffolding mechanism is deployed only for problems that are truly beyond the model’s current reach. This strategic patience fosters more resilient and independent problem-solving abilities, leading to superior overall performance.